### v1

# Point-MoE: Towards Cross-Domain Generalization in 3D Semantic Segmentation via Mixture-of-Experts

Abstract
While scaling laws have transformed natural language processing and computer vision, 3D point cloud understanding has yet to reach that stage. This can be attributed to both the comparatively smaller scale of 3D datasets, as well as the disparate sources of the data itself. Point clouds are captured by diverse sensors (e.g., depth cameras, LiDAR) across varied domains (e.g., indoor, outdoor), each introducing unique scanning patterns, sampling densities, and semantic biases. Such domain heterogeneity poses a major barrier towards training unified models at scale, especially under the realistic constraint that domain labels are typically inaccessible at inference time. In this work, we propose Point-MoE, a Mixture-of-Experts architecture designed to enable large-scale, cross-domain generalization in 3D perception. We show that standard point cloud backbones degrade significantly in performance when trained on mixed-domain data, whereas Point-MoE with a simple top-$k$ routing strategy can automatically specialize experts, even without access to domain labels. Our experiments demonstrate that Point-MoE not only outperforms strong multi-domain baselines but also generalizes better to unseen domains. This work highlights a scalable path forward for 3D understanding: letting the model discover structure in diverse 3D data, rather than imposing it via manual curation or domain supervision.
1 Introduction
In both language and vision, we have witnessed remarkable advances driven by two trends: the aggregation of massive heterogeneous datasets [59, 64, 5, 57, 9, 89], and the deployment of models of ever-increasing capacity trained on them in a unified manner [8, 18, 55, 39, 25, 49]. These models succeed not because they were crafted for a specific domain, but because they were trained to learn underlying patterns across a wide variety of datasets. However, 3D point cloud understanding has yet to follow the same trajectory of progress, despite its wide applications in robotics[13, 50, 84], autonomous systems[15, 14, 44], augmented reality[90, 75], and embodied intelligence[77, 71]. Taking the 3D semantic segmentation task as an example, although diverse 3D datasets exist—such as ScanNet [21], SemanticKITTI [6], nuScenes [11], and ASE [3]—each covers only a narrow subset of the real-world diversity in 3D assets. Point clouds are generated from a variety of pipelines: RGB-D sensors [12], terrestrial and airborne LiDAR [33, 51], and multi-view stereo [83]. Each source brings its own characteristics—differing point densities, coverage patterns, reconstruction artifacts, and semantic biases (Fig. 1 left). The result is a 3D ecosystem of rich, yet siloed data sources, where models trained in one silo often fail to generalize across others.
A natural response to address this challenge is to train across such silos. For example, one might naively train the state-of-the-art Point Transformer V3 (PTv3) [78] (Fig. 1a) on mixed-domain 3D datasets. However, such models struggle to reconcile heterogeneous characteristics of the data, even with increased network capacity (Tab. 1). To mitigate this, recent methods introduce domain-aware components. Point Prompt Training (PPT) [80] (Fig. 1b) modifies PTv3 by assigning domain-specific normalization layers. Similarly, One-for-All [74] achieves improved 3D detection performance by employing a lightweight classifier to infer dataset labels, then applying domain-specific adapters. While these methods show notable gains, they rely on access to domain labels during training and exhibit performance degradation on unseen domains. A likely cause is that adapting only the normalization parameters may be insufficient to fully capture the complex variations required for scalable, cross-domain generalization.
In this work, we hypothesize that Mixture-of-Experts (MoE) architectures [35, 38] are well-suited for cross-domain joint training. We introduce Point-MoE, a sparse MoE model built upon PTv3 [78]. As illustrated in Fig. 1c, Point-MoE replaces each feed-forward layer in PTv3 with a MoE layer comprising of multiple expert networks (implemented as feed-forward MLPs) and a router that dynamically selects a sparse subset of experts for each input token. Our design is motivated by three key insights. First, MoE provides a natural inductive bias for handling diverse inputs by encouraging expert specialization and learning token-to-expert routing [17, 82]. This allows a single model to be trained on mixed-domain data without requiring access to domain labels. Second, compared to normalization-based domain adaptation methods like PPT [80] and One-For-All [74], MoE offers a significantly larger and more expressive representation space, enabling better modeling of domain-specific variations. Third, sparse MoE architectures improve computational efficiency during both training and inference, and have emerged as a standard design for scaling large models [41, 28, 26, 58, 22]. We extend these empirical benefits of MoE, previously established in language and vision, to the domain of 3D point cloud understanding, with a particular focus on 3D semantic segmentation.
Figure 1: Overview of cross-domain training architectures. Point clouds exhibit diverse characteristics across domains. (a) Naively training Point Transformer V3 (PTv3) [78] on multi-domain data leads to degraded performance within each domain. (b) Point Prompt Training (PPT) [80] addresses this by adding domain-aware normalization parameters. (c) Our proposed Point-MoE tackles this challenge with Mixture-of-Experts (MoE), enabling dynamic expert specialization across domains. Domain labels can optionally be provided to the router to guide the expert selection.
We conduct extensive experiments across a diverse mixture of 3D datasets, spanning indoor and outdoor scenes, synthetic and real data, and multiple sensor types. To comprehensively assess the effectiveness of Point-MoE, we evaluate it under both domain-agnostic (i.e., without access to domain labels) and domain-aware training regimes. Our results show that Point-MoE consistently adapts well to both settings, achieving strong generalization to unseen domains while outperforming competitive baselines, with improved training (Fig. 2) and inference efficiency (Tab. 1 and 2). We also conduct a thorough ablation study analyzing key design choices, including the number of experts, sparsity level, and the placement of MoE layers (Tab. 3). Remarkably, even without domain supervision, Point-MoE learns to route input data meaningfully and to specialize experts organically—emerging from the data distribution itself (Fig. 3 and 4).
To the best of our knowledge, this work presents the first systematic exploration of the Mixture-of-Experts architecture for 3D point cloud understanding in the context of cross-domain generalization. These findings point toward a scalable path forward for 3D perception: instead of designing separate models for each dataset or domain, we should aim to develop unified systems that can adapt across the full spectrum of 3D data sources. In this regard, Point-MoE emerges as a promising and flexible framework for bridging domain gaps in large-scale 3D semantic segmentation.
2 Related Work
Point Cloud understanding.
Deep learning backbones for 3D scene understanding can be broadly categorized by how they process point clouds: projection-based methods [15, 40, 42, 62], voxel-based methods [48, 61, 19, 29], and point-based methods [52, 53, 67, 86]. More recently, transformer-based architectures operating on point clouds have emerged as strong contenders in this space [31, 79, 87, 78]. Among them, Point Transformer V3 (PTv3) [78] represents the current state of the art, introducing a novel space-filling curve based serialization scheme, that transforms the unstructured 3D points into a 1D sequence, enabling efficient transformer-based modeling. A recent series of works have explored scaling such point cloud models to the multi-dataset setting using domain-aware training strategies. Point Prompt Training (PPT)  [80] demonstrates that using domain-specific normalization layers during pretraining improves 3D segmentation performance across multiple datasets. In 3D detection, One-for-All [74] adopts a sparse convolutional network with domain-specific partitions and normalization layers to boost multi-domain performance. Rather than relying on such domain-specific architectural branches [80, 74], we propose a unified sparse expert framework that learns to specialize dynamically across various domains, without requiring explicit dataset labels during training or inference.
Mixture-of-Experts.
Mixture-of-Experts (MoE) [35, 38] aims to increase model capacity without a proportionate increase in compute costs by levereaging conditional computation – large portions of the model are only activated conditioned on certain input samples [60]. In NLP, MoE has enabled scaling up models from hundreds of millions to hundreds of billions (and beyond) parameters: GShard scales to 600 B parameters [41], Switch Transformer to 1.6 T parameters [28], ST-MoE to 260 B parameters [91], and GLaM to 1.2 T parameters [26], all at roughly the same training cost as a dense model one-tenth the size. V-MoE [58] applies this recipe to vision, demonstrating competitive accuracy at 15 B parameters with batch-priority routing. More recently, Mixtral 8x7B [37] and DeepSeek models [22, 23, 24] all employ Mixture-of-Experts for large language model pretraining.
Apart from being a powerful tool for scaling up model capacity, MoE is well-suited to training on data drawn from heterogeneous domains, because the experts naturally tend to self-organize on a per-domain basis. Early work shows gains in multi-source sentiment and POS tagging [30], domain-adaptive language modeling [32], automatic chart understanding [82], and multilingual translation [17]. Mod-Squad [16] introduces a modular mixture-of-experts framework for multi-task learning, optimizing task-expert matching to balance cooperation and specialization across tasks. DAMEX [36] applies this to visual domains, while MoE-LLaVA [45], DeepSeek-VL [47], and DeepSeek-VL 2 [81] report similar benefits in vision-language models. To our knowledge, ours is the first attempt at extending MoE to address multi-dataset training and cross-domain generalization for 3D point clouds.
3 Point-MoE for Cross-Domain Point Cloud Understanding
Task definition. Given a 3D scene represented as a point cloud $\mathcal{P}=\{p_{i}\}_{i=1}^{n}$, where each point $p_{i}\in\mathbb{R}^{3}$ denotes coordinates, semantic segmentation aims to predict a single class label $\hat{y}_{i}\in\mathcal{C}$ to each point $p_{i}$ from a fixed label set $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{m}\}$. Multi-domain data are utilized during training. Let the domain containing $N$ point clouds be $\mathcal{D}=\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left(p_{i},y_{i}}\right)_{i=1}^{n}}\right\}^{N}$. Then the complete training set across $d$ domains can be represented as: $\mathopen{}\mathclose{{}\left\{\mathcal{D}_{k}}\right\}_{k=1}^{d}$. The goal of multi-domain joint training is to learn a unified model over all domains that minimizes prediction error across $\mathopen{}\mathclose{{}\left\{\mathcal{D}_{k}}\right\}$, while also generalizing effectively to domains unseen during training.
We adopt a minimal Mixture-of-Experts design that closely follows the standard architecture prevalent in the recent NLP literature. This simplicity allows us to leverage existing scalable Transformer-based MoE implementations with minimal modification. For the base model, we use Point Transformer V3 (PTv3) [78], a state-of-the-art architecture for point cloud understanding. Below we introduce the architecture of Point-MoE in detail.
Mixture-of-Experts layer. The core of MoE is to route input tokens dynamically to specialized subnetworks, referred to as experts, via a lightweight gating mechanism. Given an input feature vector $\mathbf{x}\in\mathbb{R}^{d}$, each MoE layer contains $N$ experts $\{f_{i}\}_{i=1}^{N}$ and a gating network $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N}$. For each input $\mathbf{x}$, the MoE layer output is computed as a weighted sum over the top-$k$ activated experts:
$\mathrm{MoE}(\mathbf{x})=\sum_{i\in\mathcal{S}_{\mathbf{x}}}G_{i}(\mathbf{x})f_{i}(\mathbf{x})$ (1)
where $\mathcal{S}_{\mathbf{x}}$ denotes the set of top-$k$ experts selected for input $\mathbf{x}$, and $G_{i}(\mathbf{x})$ represents the normalized gating weight assigned to expert $i$. In practice, the gating function is typically implemented via a linear projection followed by a sparse softmax, ensuring activation of only a small subset of experts, with $k\ll N$. Although auxiliary losses for balancing expert load are common in prior work [60], we empirically observed them to be unnecessary and therefore excluded them from our final setup.
Integration into PTv3. To tailor the MoE architecture specifically for 3D point cloud data, we integrate the MoE mechanism directly into PTv3 by replacing the shared linear projection layers in PTv3’s attention module with an MoE layer. The original linear projections mapping input features into key, query, and value vectors within the attention mechanism are replaced by MoE layers, while all subsequent computations remain unchanged. This integration allows the MoE experts to specialize naturally in domain-specific transformations of point cloud data. Fig. 1c provides a visual overview of the modified architecture. It is crucial to insert the MoE layer before the post-normalization; otherwise, we observe a significant reduction in performance gains.
Language-guided classification. To bridge label discrepancies across datasets, we project per-point features to a shared language space using CLIP [56] text embeddings, following PPT [80]. This enables supervision via class names, without relying on dataset labels. For instance, the class “pillow” is missing in ScanNet (grouped as “other”) but is explicitly labeled in Structured3D. Our model aligns point features with the CLIP [56] embedding of “pillow,” allowing knowledge to transfer across domains even in the absence of direct supervision.
Domain-aware gating. Our preliminary experiments show that while basic Point-MoE naturally specializes experts implicitly, providing explicit domain information could further guide and accelerate expert specialization. To explore this, we introduce a dataset-aware gating variant termed Domain-Aware Point-MoE. In this variant, we explicitly associate each training domain $\mathcal{D}_{i}$ with a unique, learnable domain embedding vector $d_{i}\in\mathbb{R}^{d}$. Before routed through the gating mechanism, each input feature $x_{t}$ is concatenated with the corresponding domain embedding: $G_{t}(d)=\mathrm{softmax}(W\cdot(x_{t}\oplus d_{i}))$, where $W$ denotes a learnable linear projection, and the concatenation operation explicitly conditions gating decisions on the domain identity. This explicit conditioning encourages faster and more clearly delineated expert specialization aligned with dataset-specific characteristics observed during training.
Domain randomization for robust gating. While domain-aware gating allows experts to specialize based on known domain identities, this approach assumes access to explicit domain labels at inference time. In practical scenarios, such labels are often unavailable—especially when deploying models in the wild, where domain boundaries may be ambiguous or entirely unknown. While a domain detector can be employed to approximate the domain label by mapping unseen domains to the closest known ones, such detectors may introduce errors or fail to capture nuanced domain shifts. This motivates the need for a more robust mechanism that does not rely on explicit domain identification at test time. To address this, we introduce a domain randomization strategy that enables the model to generalize to unseen or mixed-domain inputs without performance degradation on seen domains.
Specifically, we define a learnable generic domain embedding and randomly assign it to 20% of the training points, replacing their true domain-specific embeddings. This forces the gating mechanism to learn domain-agnostic routing behavior, even in the presence of domain ambiguity. The remaining 80% of points retain their original domain embeddings, ensuring that the model still benefits from explicit supervision where available. In line with prior work [70, 17], we find that domain randomization improves generalization to unseen domains while preserving strong performance on those seen during training.
Training objective. We use Mixture-of-Experts (MoE) in a supervised cross-domain setting, where the goal is to train a unified model $f(x;\theta)$ across $n$ domains $\{\mathcal{D}_{i}\}_{i=1}^{n}$, each containing samples $(X,Y)\sim\mathcal{D}_{i}$. The objective is to minimize the average loss across domains:
$\min_{\theta}\sum_{i=1}^{n}\mathbb{E}_{(X,Y)\sim\mathcal{D}_{i}}\;\mathcal{L}(f(X;\theta),Y)$ (2)
We consider two batching strategies during training: (1) Single-domain batching, where each batch is sampled from only one domain $\mathcal{D}_{i}$ at a time. (2) Mixed-domain batching, where each batch contains samples from multiple domains. Our implementation is based on the official open-source code of PTv3 released by Pointcept [20].
4 Experiments
We conduct comprehensive experiments to evaluate Point-MoE for cross-domain 3D semantic segmentation. First, we describe baseline methods and experimental setups (Sec. 4.1). Then, we present detailed empirical results under two primary joint training scenarios: (1) indoor-only, and (2) combined indoor-outdoor training (Sec. 4.2). Finally, we provide ablation studies that highlight key design choices of the MoE architecture (Sec. 4.3) and detailed analysis of MoE behaviors (Sec. 4.4).
4.1 Experimental Setup
Baseline methods. Since our proposed Point-MoE framework is built upon PTv3 [78], we adopt PTv3 and PPT [80] as primary baselines. PPT, originally designed for dataset-specific BatchNorm  [34] and LayerNorm [4] layers and fine-tuning, is adapted here for joint training. Because PPT requires dataset labels at inference time, we follow One-for-All [74] to train a lightweight domain classifier (achieving over 95% accuracy), which we also use for Point-MoE under the domain-aware setting.
To ensure a fair comparison, we begin by benchmarking the dense PTv3 model under the domain-agnostic joint training setting, systematically varying key design factors such as mixed-batch composition and normalization strategy. In this setup, models are trained across multiple domains without access to dataset labels, and no label information is provided at inference time. This setting reflects a realistic deployment scenario where domain metadata is unavailable. For the domain-aware joint training setting, we compare to the PPT + One-for-All baseline. Here, models are trained with access to dataset labels. During inference, we either provide predicted labels using a trained domain classifier or assign a “generic” class to simulate deployment in unknown or mixed domains. This allows us to isolate the value of domain conditioning and assess the robustness of Point-MoE under both domain-aware and domain-agnostic regimes.
Datasets. Experiments are conducted on three indoor datasets—ScanNet [21], S3DIS [2], and Structured3D [88]—and evaluated on both seen and unseen domains, including Matterport3D [12]. We then evaluate the best-performing Point-MoE and baseline configurations under a more challenging indoor-outdoor setting, incorporating outdoor datasets from nuScenes [11], SemanticKITTI [7], and Waymo [65]. Voxel sizes are standardized to 0.02m for indoor scenes and 0.05m for outdoor scenes. Throughout, we denote model capacity using -S (small) and -L (large) suffixes for all architectures, including PTv3, Point-MoE, and PPT.
4.2 Experimental Results
Table 1: Performance of Joint Training for Indoor 3D Semantic Segmentation. All models are jointly trained on three indoor datasets and evaluated under two settings. In Domain-agnostic joint training, models are evaluated using only the point cloud and the corresponding dataset-specific label set. In Domain-aware joint training, a predicted scene label from a domain classifier is provided at inference time to guide the model. “-G” indicates that model utilizes domain randomization and contains a learnable “generic” domain label during training. All results are reported as mean Intersection-over-Union (mIoU).
Methods Params Activated Mix Layer. ScanNet Structured3D S3DIS Matterport3D
Seen Unseen
Val Val Test Val Val Test
Single-dataset training
PTv3-ScanNet [78] 46M 46M 75.0 19.8 20.9 8.8 38.9 44.5
PTv3-Structured3D [78] 46M 46M 21.2 81.5 79.4 7.1 21.0 24.0
PTv3-S3DIS [78] 46M 46M 4.9 3.5 5.3 67.6 5.0 3.4
PTv3-Matterport3D [78] 46M 46M 53.1 22.9 23.7 9.8 54.1 56.0
Domain-agnostic joint training
[…]
Domain-aware joint training
[…]
Table 2: Cross-domain 3D Semantic Segmentation across Indoor and Outdoor Scenes. We evaluate the best-performing models from the indoor-only setting in a more challenging joint training setup that spans both indoor and outdoor datasets. This setup better reflects real-world deployment scenarios, requiring models to generalize across diverse domains.
ScanNet Structured3D S3DIS nuScenes SemKITTI Matterport3D Waymo
Indoor Outdoor Indoor Outdoor
Seen Unseen
Methods Params Activated Mix Layer. Val Val Test Val Val Val Val Test Val Test
Domain-agnostic joint training
[…]
Domain-aware joint training
[…]
Single-dataset training. We begin by evaluating the standard setting in which models are trained and evaluated on individual datasets. Note we train end evaluate PTv3 with language-guided classification. As shown in Tab. 1, PTv3 achieves strong in-domain performance, but its performance drops significantly on unseen domains. For instance, a PTv3 model trained on S3DIS achieves only 3.5% on Structured3D and 4.9% on ScanNet. While domain-specific models are not expected to generalize across domains or categories, these results underscore the brittleness of single-dataset training. This motivates the need for cross-domain training strategies that expose the model to diverse data and promote broader generalization.
Domain-agnostic joint training. In the indoor-only setting, Point-MoE outperforms all variants of PTv3 in most domains and performs competitively on S3DIS. Referring to Tab. 1, Point-MoE-S achieves 75.4% on ScanNet and 69.7% on S3DIS, outperforming PTv3-S by +4.1% and +13.8%, respectively. Remarkably, Point-MoE-S activates only 52M parameters compared to 97M in PTv3-L, yet still surpasses PTv3-L by +0.8% on ScanNet and +8.4% on S3DIS. In addition, Point-MoE-L reaches 75.8% on ScanNet and 73.0% on Structured3D, significantly ahead of PTv3-L. On the unseen Matterport3D domain, Point-MoE-L obtains 41.6% on validation set and 47.8% on test set, outperforming PTv3-L by +1.4% and +1.2% respectively while activating only 59M parameters at inference. In the more challenging indoor and outdoor setting, Point-MoE-L again leads across the board with a few being competitve. Referring Tab. 2, Point-MoE-L achieves 69.2% on S3DIS validation and 70.2% on test split, outperforming the best dense baseline by +5.4% and +6.3%, respectively. On outdoor datasets, Point-MoE-L achieves 72.9% on nuScenes exceeding the dense baseline by +2.6%. Notably, on the unseen datasets, Point-MoE-L outperforming all dense PTv3 variants by a clear margin. These gains are achieved while activating nearly half the parameters of PTv3-L, underscoring the efficiency and scalability of sparse expert routing.
Domain-aware joint training. We further evaluate Point-MoE in the domain-aware setting, where models are trained with access to dataset labels. In the indoor-only setting, both domain-aware Point-MoE and its generic variant (Point-MoE-G) surpass all strong baselines: variants of PPT [80] + One-for-All [74]. As shown in Tab. 1, Point-MoE-G-L achieves 76.7% on ScanNet, 68.0% on Structured3D, and 67.0% on S3DIS, outperforming the best PPT + One-for-All baseline by +2.0%, +3.1%, and +2.3%, respectively. On the unseen Matterport3D dataset, it reaches 41.8% on the validation split and 47.8% on the test set, again outperforming all baselines. Notably, using only 20% of training data tagged as “generic” does not degrade performance on seen domains.
In the indoor-and-outdoor setting, Point-MoE-G-L continues to outperform all baselines across the board. Referring to Tab. 2, it achieves 77.0% on ScanNet, 72.1% and 74.2% on Structured3D (val/test), 62.4% on S3DIS and 72.7% on nuScenes—surpassing the best baseline by +0.3%, +4.6%, +4.1%, +1.0% and +0.5%, respectively. On the unseen domains, including Matterport3D and Waymo, Point-MoE-G-L delivers the strongest generalization. Crucially, it performs on par with domain-aware Point-MoE on seen datasets while achieving superior accuracy on unseen ones. These results validate that domain randomization with a generic label can enhance robustness without compromising in-domain performance. We still observe a performance gap between cross-domain and single-dataset training in cases like Structured3D [88]. Single-dataset models can adapt to domain-specific cues and distributions, while Point-MoE must balance capacity across domains. As a result, its performance may trail on large, homogeneous datasets, but remains competitive overall due to its generalization ability.
4.3 MoE Design Ablation
Table 3: Ablation of Point-MoE Design and Hyperparameters. We conduct a systematic ablation on key design and hyperparameter choices of Point-MoE across four indoor multi-domain datasets: ScanNet (Scan), Structured3D (S3D), S3DIS, and Matterport3D (Mat).
[…]
$10^{-4}$ 72.3 64.1 67.9 39.5
$10^{-3}$ 71.6 58.8 67.4 37.4
(a) Auxiliary load balancing loss. Removing the auxiliary loss consistently improves performance.
[…]
top 1 74.4 64.8 65.5 39.0
[…]
top 3 73.8 62.4 67.8 39.7
(b) Top-$k$ expert selection. Best performance is achieved with $k{=}2$ experts.
[…]
LN. 70.8 56.9 67.3 37.2
RMS. 45.3 36.5 51.1 24.6
(c) Normalization. BatchNorm yields strongest performance; RMSNorm degrades results.
[…]
dec. 74.0 64.9 66.6 38.1
[…]
(d) MoE position. Inserting MoE in all layers yields the best results.
[…]
(e) Activation functions. ReLU performs the best.
[…]
(f) Expert width. H indicates original feature dimension.
[…]
1 73.6 62.3 66.6 39.1
(g) Shared experts. Not sharing experts yields better performance.
[…]
(h) Number of experts. More experts improve performance.
[…]
4 74.0 64.9 66.6 38.1
[…]
(i) Batch size. Larger batchsize improves performance.
We investigate the impact of MoE design choices and hyper-parameters choices in indoor cross-domain 3D semantic segmentation. All experiments are conducted with ScanNet [21], Structured3D [88] and S3DIS [2] and evaluate without providing dataset label on seen dataset and also evaluate on unseen dataset: Matterport3D [12]. Results are presented in Tab. 3.
Auxiliary load-balancing loss. The auxiliary load-balancing loss is designed to encourage uniform expert usage and prevent expert collapse [54, 76, 43]. We ablate its effect by training models with varying strengths of this loss in Tab. 3. Interestingly, we find that removing the auxiliary loss entirely leads to consistently better performance across domains. In contrast, stronger auxiliary losses result in significant degradation. We hypothesize that this is due to the inherent imbalance in the distribution of samples across domains in 3D point cloud datasets—a phenomenon also observed by ChartMoE [82].
Effect of activated experts. Tab. 3 compares the performance of configurations with 1, 2 and 3 activated experts, finding diminishing return when increasing number of activated experts. We find that activating two experts yields the best performance across domains. This analysis suggests that merely increasing the number of activated experts does not guarantee improved performance.
MoE position. We ablate the placement of MoE layers to understand where expert specialization is most effective. Empirically, we observe greater expert diversity when MoE is applied in the decoder compared to the encoder. In Tab. 3, we compare two configurations: applying MoE layers only in the decoder, versus applying them in both the encoder and decoder. Results show that applying MoE to both stages consistently yields better performance, suggesting that expert specialization benefits from exposure to both encoder and decoder.
Effect of scaling model with MoE. Tab. 3 compares Point-MoE with 4 and 8 experts with activated experts fixed at 2. Increasing the number of experts from 4 to 8 leads to improved performance on seen domains, and yields a modest gain of 0.4% on the unseen domain. Tab. 3 ablates the expert width by varying the intermediate size of the MLP (e.g., 1H, 2H). The results show a mixed trend: wider experts improve performance on ScanNet and Matterport3D, but degrade results on Structured3D and S3DIS. Finally, we investigate the effect of adding shared experts [63, 72, 22], which are applied universally without gating. We find that disabling shared experts consistently leads to better accuracy and domain specialization across all benchmarks, as reported in Tab. 3.
Ablation of hyper-parameters choices. We provide justification for our selected hyperparameters. In Tab. 3, BatchNorm [34] consistently outperforms other normalization methods [4, 85] in the cross-domain setting. Tab. 3 shows that ReLU [1] slightly outperforms SiLU [27]. Additionally, Tab. 3 demonstrates that importance of batch size in MoE training. Specifically, the batch size must exceed the number of experts to avoid performance degradation. In general, larger batch sizes lead to improved performance across domains.
4.4 Analysis of MoE
Figure 2: Training Loss and Validation mIoU over time. Left: training loss curves for joint training across multiple domains. Right: validation mIoU on indoor datasets. Point-MoE-L achieves competitive performance to PPT, and both significantly outperform the baseline PTv3-L.
Training efficiency and validation mIoU. Fig. 2 shows the training loss and validation mIoU curves for four models: the baseline PTv3-L, its improved variant PTv3-Mix-LN (which incorporates mixed-domain batches and LayerNorm), PPT-L [80], and our proposed Point-MoE-L. All models are trained from scratch across multiple domains. Point-MoE-L converges faster and achieves strong validation mIoU without using explicit dataset labels, matching or exceeding the performance of PPT-L trained with ground-truth domain labels. While all models reach similar training loss, only Point-MoE-L and PPT-L generalize effectively—reinforcing that low training loss is not indicative of strong cross-domain performance. On ScanNet, Structured3D, Matterport3D, and nuScenes, Point-MoE-L shows consistent improvement and avoids early plateaus seen in PPT-L, especially on Structured3D, suggesting stronger long-term learning. PTv3-L fails to generalize and exhibits unstable validation curves. PTv3-Mix-LN improves stability over PTv3-L significantly, but still underperforms Point-MoE-L and PPT-L.
Figure 3: Expert Choices Visualization. Different color represents a different expert. (a) shows a specific expert focusing on edges across the scene, revealing spatially coherent routing. (b) and (c) highlight expert selections in indoor scenes, where semantically related regions (e.g., chair, desk) are consistently routed to the same expert. (d) displays an outdoor scene with sparse LiDAR points, yet the model still demonstrates meaningful routing despite limited geometry.
Expert choice visualization. Fig. 3 showcases expert assignments for a validation scene at selected layers. In (a), we observe that early encoder layers rely heavily on geometric cues for routing. For instance, green experts are consistently activated along object boundaries such as the edges of desks and chairs, while red experts dominate flat surfaces. In (b) and (c), the decoder layers exhibit more semantically meaningful expert selection—likely due to their proximity to the loss function—with distinct experts attending to objects like desks, chairs, floors, and walls. In (d), we examine an outdoor scene with sparse LiDAR data. Despite limited geometric structure, the model still organizes routing meaningfully: nearby points are routed to blue experts, while farther points activate red experts. We include more visualizations in the appendix for completeness. We also note occasional visual artifacts where isolated points are assigned different experts than their neighbors, which may be related to PTv3’s architectural choices such as point serialization or positional encoding.
Figure 4: Expert routing behavior across domains. The left panel shows the most frequently used expert paths through each encoder (E-) and decoder (D-) layer, with feature channel sizes marked in parentheses, illustrating the variation in expert selection for tokens from different datasets. The router tends to assign experts based on the input domain, particularly in the decoder MoE layers. To quantify this, we compute the Jensen-Shannon Divergence (JSD) between the expert selection distributions of different datasets at each MoE layer.
Token pathways. To understand how Point-MoE adapts to diverse domains, we analyze expert routing behavior at the token level. Specifically, we track the top-1 expert assignment for each token across all MoE layers and construct full routing trajectories from the final layer back to the first. We then identify the top 100 most frequent expert paths based on their occurrence across all tokens. As shown in Fig. 4, encoder expert paths are substantially less diverse than those in the decoder, indicating that encoder layers perform more domain-agnostic processing. Interestingly, we observe a sparse routing pattern in deeper encoder layers. This may be attributed not to feature reuse, but to the U-Net-style design where features are spatially deep but token sparsity increases, resulting in reduced variability in routing decisions.
When examining domain-level trends, we find that certain dataset pairs—such as SemanticKITTI and nuScenes or ScanNet and Structured3D—share similar expert pathways, suggesting that Point-MoE implicitly clusters domains with related geometric or semantic structures. To quantify these observations, we compute the Jensen-Shannon Divergence (JSD) [46] between expert selection distributions across datasets at each MoE layer. JSD [46] is an entropy-based measure of divergence between expert routing distributions across datasets, weighted by their token proportions; its formal definition is provided in the supplementary. As shown in Fig. 4 (right), decoder layers exhibit significantly higher JSD, indicating stronger domain-specific specialization. Several encoder layers also display nontrivial JSD, underscoring the benefit of placing MoE throughout the network.
5 Discussions
This work embraces the “bitter lesson” [66] in AI: scalable generalization emerges from flexible architectures trained on diverse data, rather than hand-engineered domain priors [68]. We introduce Point-MoE, a mixture-of-experts architecture for 3D understanding that dynamically routes tokens to specialized experts across both indoor and outdoor domains. Our experiments demonstrate that Point-MoE not only achieves competitive performance on seen datasets but, crucially, also generalizes effectively to novel domains without explicit domain labels or supervision. Through careful analysis of expert routing behavior, we show that the model self-organizes around both geometric and semantic structures, adapting its computation pathways to the nature of the input. These results highlight the potential of sparse, modular computation for robust and scalable 3D scene understanding. We provide more analysis of Point-MoE and discuss our limitations in the supplementary material.
6 Acknowledgement
The authors gratefully acknowledge Research Computing at the University of Virginia for providing the computational resources and technical support that made the results in this work possible. Zezhou Cheng acknowledges the Adobe Research Gift, as well as GPU support provided by the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program and the National Artificial Intelligence Research Resource (NAIRR) Pilot program.
Appendix
For a thorough understanding of our Point-MoE, we have compiled a detailed Appendix. The table of contents below offers a quick overview and will guide to specific sections of interest.
[…]
Appendix A Limitations, Licenses and Risks
A.1 Limitations
Our work is limited by computational resource constraints, which restrict the scale of both the Point-MoE architecture and the training datasets. While our results already demonstrate strong generalization across diverse 3D domains, larger models and a broader training corpus could further enhance performance. Additionally, our expert routing relies solely on the standard gating loss without auxiliary regularization for load balancing or diversity, which may lead to suboptimal expert specialization. Finally, the dynamic routing mechanism incurs nontrivial communication and synchronization overhead compared to dense backbones like Point Transformer V3, highlighting the need for more efficient expert parallelism or conditional computation in future work.
A.2 Artifacts and License
We report a list of licenses for all datasets and code used in our experiment in Tab. 4. We strictly follow all the datasets and code licenses and limit the scope of these datasets and code to academic research only.
Table 4: License information for the scientific artifacts used.
Data Sources URL License
ScanNet [21] Link ScanNet Terms of Use
Structured3D [88] Link Structured3D Terms of Use
S3DIS [2] Link CC BY-NC 4.0
Matterport3D [12] Link Matterport3D License
nuScenes [10] Link CC BY-NC-SA 4.0
SemanticKITTI [6] Link CC BY-NC-SA 3.0
Waymo Open Dataset [73] Link Waymo Research License
Software Code URL License
Pointcept [20] Link MIT License
A.3 Ethical concerns and Risks
This study does not involve human subjects or annotators. All experiments are performed on publicly released 3D benchmarks—ScanNet [21], Structured3D [88], S3DIS [2], Matterport3D [12], SemanticKITTI [6], nuScenes [10], and Waymo [65]—each distributed under their respective licences. Although these datasets are widely used, they may still reflect geographic, socioeconomic, or sensor–specific biases, which Point-MoE could inherit or amplify when deployed in real-world applications. We therefore encourage future work to develop debiasing techniques, to curate more balanced point-cloud corpora, and to incorporate fairness-oriented evaluation metrics so that cross-domain 3D models can serve all communities responsibly.
Appendix B Additional Experiments
B.1 Experiment Settings
Data. We compare a diverse set of 3D semantic segmentation datasets spanning both indoor and outdoor environments. ScanNet [21] is a large-scale indoor RGB-D dataset reconstructed from real-world scans of apartments and offices. S3DIS [2] contains real indoor scans of commercial buildings, providing richly annotated scenes with diverse architectural structures. Structured3D [88] is a synthetic dataset based on photo-realistic rendered scenes from CAD models, offering clean labels and dense annotations. Matterport3D [12] consists of real RGB-D scans captured in a wide range of residential environments, covering multiple floors and room types.
For outdoor settings, SemanticKITTI [6] provides point clouds captured by a 64-beam LiDAR mounted on a car, annotated for semantic understanding of dynamic street scenes. Waymo Open Dataset [73] extends this setup with more diverse driving environments and higher annotation quality, collected using a custom autonomous vehicle sensor suite. nuScenes [10] offers multimodal driving data, including LiDAR scans, captured in urban environments with complex traffic scenarios and longer temporal contexts. We include detailed dataset statistic in Tab. 5.
Furthermore, we provide detailed per-class comparisons on semantic classes for both indoor and outdoor datasets, as shown in Fig. 8 and Fig. 8.
Table 5: Overview of 3D Semantic Segmentation datasets differences. L, W, and H represent the length, width, and height. For all datasets, we use the raw point clouds to obtain the point ranges.
Dataset Sensor Point Range (m) Scene Type
ScanNet [21] Reconstructed L=[-2.13, 17.12], W=[-1.46, 18.19], H=[-0.38, 6.97] indoor
S3DIS [2] Reconstructed L=[-37.93, 29.93], W=[-26.08, 46.06], H=[-2.65, 6.58] indoor
Structured3D [88] Synthetic L=[-53.70, 52.05], W=[-82.07, 53.43], H=[-0.78, 15.49] indoor
Matterport3D [12] RGB-D Camera L=[-34.99, 72.17], W=[-57.85, 75.65], H=[-6.79, 10.21] indoor
SemanticKITTI [6] 64-beam LiDAR L=[-81.15, 80.00], W=[-80.00, 79.99], H=[-32.49, 3.51] outdoor
Waymo [73] 64-beam LiDAR L=[-74.98, 76.53], W=[-75.00, 75.00], H=[-18.36, 6.78] outdoor
nuScenes [10] 32-beam LiDAR L=[-105.17, 105.25], W=[-105.30, 105.17], H=[-53.42, 20.01] outdoor
Training. Our default joint training strategy and Point-MoE-L configs are summarized in Tab. 8. We experiment with two batching schemes: (1) single batch, where each iteration samples points exclusively from one dataset, and (2) mix batch, where samples are drawn from multiple datasets with equal probabilities. We find that larger batch sizes lead to more stable expert assignment and better convergence behavior in the Mixture-of-Experts architecture. All models are trained from scratch on the joint multi-domain datasets without any fine-tuning stage.
Table 6: Categorical settings of Indoor Datasets.
Dataset #C
wall
floor
cabinet
bed
chair
sofa
table
door
window
bookshelf
bookcase
picture
counter
desk
shelves
curtain
dresser
pillow
mirror
ceiling
refrigerator
television
shower curtain
nightstand
toilet
sink
lamp
bathtub
garbagebin
board
beam
column
clutter
otherstructure
otherfurniture
otherprop
other
Structured3D 25
ScanNet 20
S3DIS 13
Matterport3D 21
Table 7: Categorical settings of Outdoor Datasets.
Dataset #C
car
bicycle
motorcycle
truck
other-vehicle
person
bicyclist
motorcyclist
road
parking
sidewalk
other-ground
building
fence
vegetation
trunk
terrain
pole
traffic-sign
barrier
bus
construction_vehicle
pedestrian
traffic_cone
trailer
driveable_surface
other_flat
manmade
other vehicle
sign
traffic light
construction cone
tree trunk
curb
lane marker
other ground
walkable
SemanticKITTI 19
nuScenes 16
Waymo 22
Table 8: Training settings. H indicates feature dimension.
Joint-training Indoor Joint-training Indoor-Outdoor Point-MoE-L
Config Value Config Value Config Value
optimizer AdamW optimizer AdamW # experts 8
scheduler OneCycleLR scheduler OneCycleLR topk 2
learning rate 0.005 learning rate 0.005 shared expert 0
weight decay 0.05 weight decay 0.05 activation func. ReLU
batch size 6 batch size 16 auxiliary loss 0
datasets ScanNet (1), S3DIS (1), Struct.3D (1) datasets ScanNet (1), S3DIS (1), Struct.3D (1), nuScenes (1), SemanticKITTI (1) normalization BatchNorm
domain-aware False
domain rand. False
iters 120k iters 160k expert capacity 2H
(a) Detector trained on indoor domains
(b) Detector trained on indoor-outdoor domains
Figure 5: Domain Classification Confusion Matrices. (a) Confusion matrix for a domain detector trained only on indoor datasets: Structured3D [88], ScanNet [21], and S3DIS [2]. (b) Confusion matrix for a detector trained on both indoor and outdoor datasets: Structured3D [88], ScanNet [21], S3DIS [2], SemanticKITTI [6], and nuScenes [10]. In both cases, we evaluate generalization to unseen domains. Values are normalized per predicted domain (row-wise).
B.2 Additional Results and Analysis
Domain Detector. To enable domain-aware models at inference time—where ground-truth dataset labels are unavailable—we train a lightweight classifier to predict the domain of an input point cloud. This domain detector serves a purely functional role: it allows domain-conditioned variants of Point-MoE and other baselines to operate in realistic deployment scenarios where users may supply point clouds from unknown or mixed sources. During training, we observe that strong data augmentations (e.g., grid dropout or random cropping) reduce the classifier’s reliability by masking domain-specific cues. To preserve dataset-level characteristics critical for this task, we disable most augmentations and apply only minimal point cropping for efficiency. The domain detector converges quickly, with accuracy stabilizing after just a few epochs. As shown in Fig. 5, it achieves high classification accuracy—99.2% for indoor-only and 95.5% for indoor-outdoor settings—making it suitable for practical use.
The high accuracy of the domain detector suggests that 3D datasets exhibit strong dataset-specific patterns, even when covering similar semantic categories. While we do not perform a detailed analysis of the underlying factors, differences in data acquisition, such as synthetic rendering, RGB-D reconstruction, or LiDAR scanning, are likely to introduce distinct geometric and distributional biases. These biases are implicitly learnable and can be exploited by simple classifiers. For conventional models, such domain discrepancies often necessitate careful data curation or normalization strategies to improve generalization. However, our findings with Point-MoE offer a different perspective: by explicitly modeling distributional variation through expert specialization, the architecture may naturally accommodate such domain shifts without requiring aggressive data homogenization.
Does Point-MoE work on single dataset?
Table 9: Performance of Point-MoE vs. PTv3 on individual datasets. We evaluate both models independently on each dataset to examine whether Point-MoE offers benefits even without cross-domain training. Point-MoE performs comparably to PTv3.
Method ScanNet Structured3D Matterport3D nuScenes Waymo Average
PTv3 75.01 79.37 53.89 77.16 70.20 71.13
Point-MoE 75.63 79.29 53.47 78.24 69.50 71.23
We evaluate Point-MoE without cross-domain supervision by training and testing independently on three indoor datasets (ScanNet [21], Structured3D [88], Matterport3D [12]) and two outdoor datasets (nuScenes [10], Waymo [65]). On ScanNet and nuScenes, Point-MoE outperforms PTv3 by +0.62 and +1.08, respectively. On Structured3D (evaluated on its test split), performance is nearly identical (–0.08). On Matterport3D (test split) and Waymo (validation split), Point-MoE underperforms PTv3 by –0.42 and –0.70, respectively. On average, Point-MoE achieves 71.23 vs. 71.13 for PTv3 (+0.10). These results suggest that expert specialization can emerge in single-domain settings, but the overall improvement is small and not consistent across datasets. We consider the findings inconclusive.
We hypothesize that Point-MoE may require different configurations or training strategies when applied to a single dataset. In particular, the current gating and load balancing mechanisms are optimized for multi-domain settings, where diverse token distributions benefit from expert diversity. In contrast, single-domain training may benefit from fewer experts, a larger top-$k$, or stronger regularization to avoid expert under-utilization.
Table 10: Joint Training Results on Indoor 3D Semantic Segmentation without Precise Evaluator. Each model is trained jointly on three indoor datasets and evaluated under a domain-agnostic setting without relying on ground-truth domain labels during inference. We report performance when models are required to predict across all domains. To evaluate generalization with a single “generic“ label, results marked (G) use this label for both seen and unseen domains. Unmarked results rely on domain detector predictions..
Methods Params Activated Mix Batch Layer. ScanNet Structured3D S3DIS Matterport3D
Val Val Val Val
Single-dataset training
PTv3-ScanNet [78] 46M 46M 74.4 19.4 8.5 37.9
PTv3-Structured3D [78] 46M 46M 20.0 78.0 7.0 19.5
PTv3-S3DIS [78] 46M 46M 4.9 3.4 66.1 5.2
PTv3-Matterport3D [78] 46M 46M 52.7 23.0 9.5 53.6
Domain-agnostic joint training
[…]
Domain-aware joint training
[…]
Table 11: Cross-Domain 3D Semantic Segmentation on Indoor and Outdoor Scenes without Precise Evaluator. Models are trained jointly on both indoor and outdoor datasets under domain-agnostic supervision and domain-aware joint training settings. (G): Using the “generic“ label during evaluation (see definition above).
Methods Params Activated Mix Layer. ScanNet Structured3D S3DIS nuScenes SemKITTI Matterport3D Waymo
Val Val Val Val Val Val Val
Domain-agnostic joint training
[…]
Domain-aware joint training
[…]
Table 12: Joint Training Performance on Indoor 3D Semantic Segmentation with Ground-Truth Domain Labels. We additionally evaluate models trained on indoor datasets using domain-aware supervision, where explicit ground-truth domain labels are provided during training to guide the learning process.
Methods Params Activated Mix Batch ScanNet Structured3D S3DIS
Val Val Test Val
Domain-aware joint training with G.T. labels
PPT-S [80, 74] 47M 47M 74.7 60.9 65.5 71.6
PPT-L [80, 74] 98M 98M 74.6 60.9 62.6 69.2
Point-MoE-S 59M 52M 76.7 66.7 72.3 71.9
Point-MoE-L 100M 60M 75.6 69.7 71.3 69.6
Point-MoE-G-S 59M 52M 75.7 69.1 70.8 68.1
Point-MoE-G-L 100M 60M 76.7 68.1 69.6 67.1
Table 13: Joint Training Performance on Indoor-Outdoor 3D Semantic Segmentation with Ground-Truth Domain Labels. We evaluate models trained on indoor-outdoor datasets using domain-aware supervision with ground truth labels.
Methods Params Activated Mix ScanNet Structured3D S3DIS nuScenes SemKITTI
Val Val Test Val Val Val
Domain-aware joint training with G.T. labels
PPT-L [80, 74] 98M 98M 76.7 67.2 66.0 71.2 72.6 66.9
Point-MoE-L 100M 60M 76.8 69.6 70.1 70.6 72.4 64.5
Point-MoE-G-L 100M 60M 77.0 74.9 73.2 71.6 73.0 64.3
B.3 Additional Main Results
Results without precise evaluator. The results reported in Tab. 1 and Tab. 2 are obtained using the precise evaluator provided by Pointcept [20]. The precise evaluator first splits each scene into multiple overlapping chunks, then performs model inference on each chunk individually. After inference, it aggregates predictions across chunks through a voting mechanism to refine the final output. We typically observe a gain of 2–4% in accuracy on each dataset using this method. While effective, the precise evaluator introduces additional runtime overhead during evaluation. Moreover, we note that this form of ensembling may hurt overall performance when the model’s predictions are unreliable in early chunks, as incorrect outputs are also accumulated in the final vote. To support convenient and reproducible benchmarking for the community, we include results with the precise evaluator in the main text and provide additional baseline results without it elsewhere in the paper.
From both Tab. 1 and Tab. 2, we observe that Point-MoE outperforms the baselines in most domains in both the indoor-only and indoor-outdoor joint training settings. These results are consistent with those obtained using the precise evaluator reported in the main text. We do not observe any contradictory trends, as the precise evaluator typically improves all methods by a similar margin of 2–4%. As such, our analysis and conclusions align closely with those presented in the main text. How Strong is Domain-Aware Point-MoE? In the main text, we compare domain-agnostic and domain-aware settings and observe that domain-aware Point-MoE does not consistently outperform its domain-agnostic counterpart. Here we want to offer discussions and analysis to explain the reason. One contributing factor is the reliance on the domain predictor as we detailed in sec. B.2. Although the domain predictor achieves high classification accuracy, even minor errors can lead to noticeable performance degradation—either by misrouting inputs to suboptimal experts in Point-MoE or by selecting incorrect normalization statistics in PPT. To better isolate this effect, we include results in Tab. 12 and Tab. 13, where we evaluate domain-aware Point-MoE using ground-truth domain labels.
In general, Point-MoE continues to outperform standard baselines in both settings. Moreover, we find that domain-aware Point-MoE with access to ground-truth labels achieves better performance on seen domains, while domain-agnostic Point-MoE trained with domain-randomized data generalizes better to unseen domains. Interestingly, we also observe that Point-MoE trained with generic (i.e., domain-agnostic) supervision performs well even on seen domains when exposed to more diverse training inputs. This suggests that domain-randomization during training may benefit performance across both seen and unseen domains in challenging settings. Finally, we note that domain-aware Point-MoE can yield stronger results when accurate domain labels are available at inference time.
Appendix C Additional Visualization and Discussions on Point-MoE
C.1 t-SNE
Figure 6: t-SNE Visualization. The first three columns illustrate feature similarities and differences across PTv3, PPT, and Point-MoE for both encoder and decoder layers. The rightmost column highlights how Point-MoE generalizes to unseen domains, with Matterport3D features aligning closely with the most relevant seen domains—demonstrating robust cross-domain generalization.
Fig. 6 presents t-SNE visualizations [69] of encoder and decoder features from PTv3, PPT [80], and Point-MoE. PTv3 exhibits no clear separation in either encoder or decoder spaces, indicating its inability to structure domain-specific representations. In contrast, PPT shows strongly separated clusters for each domain at both encoder and decoder levels, suggesting it overfits to domain-specific patterns. While this benefits seen domains, it may hinder generalization to unseen settings.
Point-MoE, however, reveals a more interesting behavior. Its encoder features remain entangled across domains—indicating shared representation learning—while its decoder learns to disentangle domain-specific structures. This division of labor suggests that Point-MoE effectively performs domain inference within the model, without requiring explicit domain labels. Most notably, in the unseen Matterport3D domain, Point-MoE maps samples to nearby clusters of the most relevant seen domains (e.g., ScanNet), enabling strong generalization. This implicit alignment underscores the model’s ability to organize and interpret new environments by leveraging expertise learned from diverse sources.
C.2 Semantic Concept Distribution over Experts
Expert Specialization Word Cloud. To visualize expert specialization, we collect the semantic predictions from each expert in the decoder layers of Point-MoE. During inference, we log which expert is selected for each point and record the corresponding semantic prediction. For each expert, we compute the frequency distribution of predicted classes across the entire validation set and display the most frequent class as a word cloud. Colors indicate the originating dataset of the prediction (e.g., ScanNet, S3DIS, etc.), providing insight into cross-domain consistency.
We observe several interesting patterns in the expert specialization behavior. (1) In each decoder layer, there is typically at least one expert with a strong preference for outdoor data. For example, Decoder 0_0 Expert 7, Decoder 1_1 Expert 1, and Decoder 2_1 Expert 2 predominantly handle outdoor scenes. (2) Generic or ambiguous semantic classes such as otherfurniture or other-ground never dominate the assignment of any expert, suggesting that these classes are more diffusely distributed and do not concentrate through the routing mechanism.
From the visualization in Fig. 7, we observe several notable trends. Specialization emerges clearly: within each decoder block, experts tend to focus on semantically coherent regions such as bed, bathtub, window, or trailer, suggesting that experts develop functional roles under shared supervision. In many blocks, multiple experts cover similar classes, yet differ in dataset association, indicating potential cross-domain generalization. For instance, certain experts across blocks frequently predict bed or toilet across multiple indoor datasets, hinting at robust feature preferences. We also observe signs of dataset sensitivity. Classes like trailer, parking, and terrain appear predominantly in SemanticKITTI and nuScenes, while sink, shower curtain, and dresser are more common in Structured3D. Comparing across decoder layers, earlier blocks display a broader mix of classes per expert, while deeper blocks exhibit more focused and object-specific specialization, reflecting an increasing semantic refinement through the decoding process.
These findings suggest that Point-MoE encourages structured representation learning, where each expert in a given layer implicitly captures recurring semantic or geometric patterns, even without explicit specialization objectives or layer-to-layer consistency.
Figure 7: Expert Specialization Word Cloud. All classes’ frequencies of selecting their top-1 expert are measured on the validation set. Colors indicate the originating dataset of the class (e.g., ScanNet, S3DIS, etc.), providing insight into cross-domain consistency.
Expert Probability Stack Visualization. To complement the word cloud analysis, we provide fine-grained views of expert behavior in Fig. 8, Fig. 9, Fig. 10 and Fig. 11. For each semantic class, we visualize the soft routing probabilities across all experts within a specific decoder block. Each horizontal bar corresponds to a class, segmented proportionally by the probability of top-1 expert assignment to each expert. This representation allows us to examine how responsibilities are distributed across experts for different categories.
We observe that certain semantic classes are strongly associated with a small number of experts. For example, bed, bathtub, and toilet are consistently routed to a narrow subset of experts, reflecting high specialization. In contrast, classes such as wall and vegetation are distributed across multiple experts, likely due to their spatial extent and ubiquity in scenes. Some experts show broad activation across diverse classes, while others are more selective, indicating a balance between general-purpose and task-specific routing. Several consistent patterns emerge in the decoder. In Decoder Layer 3 Block 0, both truck and trailer are primarily assigned to Expert 3, suggesting structural similarity. In Decoder Layer 3 Block 1, dresser, nightstand, and counter are all routed to Expert 2, reflecting shared visual or functional features. In Decoder Layer 0 Block 1, traffic cone and barrier exhibit nearly identical routing distributions, likely due to their geometric similarity and outdoor context. These observations confirm that the model learns structured yet flexible routing strategies, shaped by both semantic identity and spatial context.
Figure 8: Expert Choices Visualization for Decoder Layer 0.
Figure 9: Expert Choices Visualization for Decoder Layer 1.
Figure 10: Expert Choices Visualization for Decoder Layer 2.
Figure 11: Expert Choices Visualization for Decoder Layer 3.
C.3 Expert Evolution.
Figure 12: JSD Dynamics During Training. We visualize the Jensen-Shannon Divergence (JSD) between expert assignment distributions across datasets to study how expert specialization emerges over time. The plots illustrate JSD dynamics across training iterations for both encoder and decoder layers.
In Fig. 12, we track the Jensen-Shannon Divergence (JSD) between expert assignment distributions across different datasets every 10k training iterations to analyze how expert specialization evolves over time. Let $P_{1},P_{2},\dots,P_{n}$ denote the expert routing distributions for each of the $n$ datasets, where each $P_{i}$ is a categorical distribution over experts. Specifically, let $P_{i}(e)$ denote the fraction of tokens from dataset $i$ that are routed to expert $e$. Let $\pi_{1},\pi_{2},\dots,\pi_{n}$ be the corresponding weights assigned to each dataset, satisfying $\sum_{i=1}^{n}\pi_{i}=1$. These weights reflect the proportion of tokens contributed by each dataset. The Jensen-Shannon Divergence is defined as:
$\text{JSD}_{\pi}(P_{1},P_{2},\dots,P_{n})=H\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}\pi_{i}P_{i}}\right)-\sum_{i=1}^{n}\pi_{i}H(P_{i}),$
where $H(P)=-\sum_{x}P(x)\log P(x)$ denotes the Shannon entropy. This formulation captures the degree of divergence between datasets in terms of how they utilize the experts, providing a measure of specialization and separation over time.
In the early stages (10k–30k), JSD values remain low across most experts, indicating that the routing behavior is largely uniform across datasets and experts have not yet developed distinct roles. Between 40k and 60k iterations, we begin to observe divergence among certain experts, marking the onset of specialization. This trend continues and becomes more pronounced by 80k iterations, where several experts display increasingly distinct assignment distributions. By 100k iterations, the JSD curves stabilize, suggesting that expert roles have converged. Some experts diverge earlier and more sharply, indicating faster specialization, while others evolve more gradually. Overall, this analysis confirms that expert diversity in Point-MoE is not predefined but emerges progressively through training.
In the encoder, we observe that some experts maintain JSD values close to zero throughout training, suggesting that their routing behavior remains largely random and undifferentiated. This observation aligns with Fig. 4, where certain encoder experts do not exhibit clear dataset- or task-specific specialization. This raises the possibility that not all encoder layers require MoE modules, or that additional regularization or training signals may be needed to encourage expert differentiation in the encoder. Nonetheless, we do observe that some encoder experts do specialize over time, indicating that meaningful expert roles can still emerge under the right conditions.
In the decoder, we observe that all experts exhibit high JSD values, indicating that their routing behavior becomes highly specialized and dataset-dependent. Additional visualizations of expert assignments can be found in Sec. C.4, further illustrating this specialization.
C.4 More Expert Choice Visualization
Fig. 13, Fig. 14, Fig. 15, Fig. 16, Fig. 17, Fig. 18, and Fig. 19 showcase expert assignments for a validation scene across all layers. Due to the hierarchical structure of the PTv3 backbone, we observe denser point cloud representations at the input and output stages, with sparser intermediate features. This design leads to visibly varying point densities and detail levels across the network. We also observe block-like artifacts in some visualizations, which are likely attributed to PTv3’s point serialization strategy. In the encoder, certain experts occasionally exhibit collapsed or spatially clustered behavior—a phenomenon we refer to as “expert clasping.” A deeper investigation into this behavior is left for future work. In the decoder layers, expert assignments closely resemble semantic segmentation outputs, even though the visualizations reflect only expert routing decisions rather than predicted labels. This suggests that experts naturally specialize in semantically coherent regions, despite the absence of explicit supervision guiding their roles.
In Fig. 13, we observe that early encoder layers primarily rely on geometric cues for expert routing. For example, certain experts—such as the green one—are consistently activated along object boundaries, including the edges of desks and chairs, while others tend to dominate on flat or horizontal surfaces. In the decoder layers, expert assignments become more semantically meaningful, with different experts focusing on specific object categories such as desks, chairs, walls, and floors.
In Fig. 14, we observe a similar overall pattern to ScanNet [21]. In the encoder, expert assignments are primarily driven by local geometric structure, with clear attention to furniture boundaries and wall intersections. Notably, edge-aware routing persists deeper into the encoder—for example, in encoder layer 2 block 0, certain experts remain focused on object contours. In the decoder, expert activation becomes increasingly aligned with semantic regions, such as tables, chairs, and structural elements. Despite the architectural and layout differences from ScanNet, the expert routing in S3DIS remains semantically consistent across the scene.
Fig. 15 illustrates expert routing on Structured3D [88], a synthetic dataset. Compared to real-world datasets, the expert assignment patterns appear sharper and more consistent, likely due to the absence of sensor noise and reconstruction artifacts. Some experts are clearly focused on flat regions, such as beds or floor. In the decoder layers, expert specialization aligns closely with object instances, and the separation across semantic regions is more distinct and well-defined, highlighting the benefits of clean, synthetic geometry for expert differentiation.
Fig. 16 presents expert routing on Matterport3D [12], which contains complex residential environments. In the early encoder layers, we observe clear expert usage patterns—for instance, a distinct separation between vertical and horizontal surfaces. In the decoder layers, expert assignments become more semantically meaningful, with experts specializing in large structural elements such as floors and ceilings, as well as common household objects like counters, sofas, and chairs. Notably, Matterport3D is not included in the training set, so this behavior reflects Point-MoE’s ability to generalize expert specialization to unseen domains.
In Fig. 17, we examine an outdoor scene from SemanticKITTI [6], characterized by sparse LiDAR data. Despite the limited geometric density, the model still exhibits meaningful expert routing: nearby points tend to be assigned to certain experts, while distant points activate others. Although the transition from geometry-driven to semantically informed expert assignment remains observable, it appears noisier due to the inherent sparsity and irregularity of the input point cloud.
Fig. 18 shows expert routing on nuScenes [10], an outdoor dataset with lower-resolution LiDAR data. In the encoder layers, expert assignments appear more fragmented, likely due to the reduced point density. Nonetheless, we observe that closer points are often routed to specific experts, while distant points are handled by others, indicating a spatially aware routing pattern. In the decoder layers, expert assignments become more semantically coherent, with specialization emerging for vehicle bodies, building facades, and road surfaces based on point position and distance.
Fig. 19 presents expert routing on Waymo [65], an unseen dataset characterized by 64-beam LiDAR scans. In the encoder, experts exhibit structured activation patterns even in early layers, with clear delineation between flat surfaces and protruding objects. In the decoder layers, expert routing becomes more semantically aligned, with experts consistently activated over road surfaces, vehicle bodies, and trees. The assigned regions remain compact and well-separated.
Figure 13: Point-MoE’s Expert Choice Visualization on ScanNet. Each point is colored by its assigned expert across encoder and decoder blocks. Early layers show structured expert usage, while deeper layers exhibit more mixed routing, reflecting the evolving representation dynamics throughout the network.
Figure 14: Point-MoE’s Expert Choice Visualization on S3DIS. Each point is color-coded based on the expert it was routed to across both encoder and decoder blocks. The earlier layers demonstrate clear structure in expert assignments, whereas the later layers reveal more diverse routing patterns, indicating evolving feature representations.
Figure 15: Point-MoE’s Expert Choice Visualization on Structured3D. Points are colored by their designated experts in the encoder and decoder. While the initial layers display consistent expert allocation, the deeper layers show a blend of routing decisions, highlighting the network’s shifting representational strategy.
Figure 16: Point-MoE’s Expert Choice Visualization on Matterport3D. Color represents the expert assigned to each point throughout the encoder and decoder. Structured expert usage is visible in early stages, transitioning to more varied routing in deeper layers, signaling changes in how features are processed.
Figure 17: Point-MoE’s Expert Choice Visualization on SemanticKITTI. The expert responsible for each point, shown through color, spans the encoder and decoder blocks. Initially, expert usage is highly organized, but becomes more dispersed in later layers, capturing the transformation in representation over depth.
Figure 18: Point-MoE’s Expert Choice Visualization on NuScene. Each point is colored according to its corresponding expert across the network’s layers. Early encoder and decoder blocks exhibit distinct expert patterns, while deeper layers demonstrate a more distributed and dynamic routing behavior.
Figure 19: Point-MoE’s Expert Choice Visualization on Waymo. Color indicates the expert each point is assigned to across encoder and decoder stages. Early network layers maintain structured expert routes, but this structure relaxes in deeper layers, illustrating evolving internal representations.